8. An investor just bought a two-year zero-coupon bond of nominal value 100$ with yield 7%. After one year, he sold it with a yield 5%. What is the rate investor? of profit gained by the
Answer #1
Price when bought = 100/1.07^ 2 = $         87.34
Sale price = 100/1.05 = $         95.24
Rate of profit = 95.24/87.34 - 1 = 9.05%